OpenStudy (anonymous):

I need help with the Pythagorean Therom my triangles sides are both 8

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you want the hypotenuse right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

then multiply the length of the side by \(\sqrt2\)

OpenStudy (misty1212):

that is always the case if the sides are the same let's call them \(a\) then by pythgoras \[a^2+a^2=h^2\\ 2a^2=h^2\\ \sqrt{2a^2}=h\\ a\sqrt2=h\]

OpenStudy (misty1212):

in your example \(a=8\) so \(h=8\sqrt2\)

OpenStudy (anonymous):

Thank you